4-23 RATIO ANALYSIS Data for Barry Computer Co. and its industry averages follow. The firms debt is priced at par, so the maPrepare a performance report on Barry Computer Company. (Problem 4-23 on pages 140-141 of the course text provides a balance sheet and an income statement for the company.)

  • Prepare your performance report to show calculations for the 14 ratios listed on page 141, as well as a comparison of your computed ratios with the listed industry averages.
  • Write a short memo to your supervisor explaining your findings and your recommendations for improvement.
  • Suggest some ways in which the company can plan to improve below industry average ratio performance.
  • Explain why your recommendations would be effective.
  • Be sure to list your computations in an appendix to your report.
Financial Ratios Formula
Current Current assets/Current liabilities
Quick Current assets-inventory/Current liabilities
Days Sales Outstanding Accounts receivable/Sales per day
Inventory Turnover Cost of goods sold/Inventories
Total Assets Turnover Sales/Total assets
Profit Margin Net income/Sales
ROA Net income/Total assets
ROE Net income/Equity
ROIC EBIT(1-T)/Total capital
TIE EBIT/Interest charges
Debt/Total Capital Debt capital/Total Capital
M/B Market price per share/Book value per share
P/E Price per share/Earnings per share
EV/EBITDA Enterprise Value/EBITDA
Answer #1

Financial Ratios

Formula

Ratios Fig.

Current

Current assets/Current liabilities

1.98x

Quick

Current assets-inventory/Current liabilities

1.25x

Days Sales Outstanding

Accounts receivable/Sales per day

76 days

Inventory Turnover

Cost of goods sold/Inventories

5.13x

Total Assets Turnover

Sales/Total assets

1.70x

Profit Margin

Net income/Sales

1.7%

ROA

Net income/Total assets

2.88%

ROE

Net income/Equity

7.56%

ROIC

EBIT(1-T)/Total capital

6.80%

TIE

EBIT/Interest charges

2.86x

Debt/Total Capital

Debt capital/Total Capital

41.54%

M/B

Market price per share/Book value per share

1.2

P/E

Price per share/Earnings per share

15.87

EV/EBITDA

Enterprise Value/EBITDA

4.84

ANALYSIS:

After evaluating the ratios, it can be observed that on one side – a few ratios are almost up-to-mark and thus nothing is to be done in that regard. But, on the other hand, a few aspects are very poor and need attention of the management or else the company could land in a crisis.

The major focus points are:

  • Day sales outstanding has a whooping 76 days credit period which will create a cash-crunch in the company due to no-rotation of funds.
  • Inventory Turnover should also be looked over. The sales team is not able to match the current level of production.
  • Total assets turnover, ROA, ROE depict that the assets and owner’s equity are not being adequately utilised. Both Sales and net income are way behind the industrial avg. parameters in terms of these ratios.
  • M/B and P/E ratios depict that the market value of company’s share is not satisfactory. The reason may its reputation in the market due to the above-mentioned deficiencies.
  • The Enterprise value (i.e., total capital – cash & cash equivalents) is very less as compared to the industrial average.

RECOMMENDATIONS:

The management must focus on its internal control, i.e.,

  • Cost management of product
  • Production efficiency
  • lead time of procurement & distribution channels
  • Credit collection process must be improved

Also, the company has to perform well and maintain a high goodwill in order to raise its market price per share.

APPENDIX – CALCULATIONS:

I have only mentioned a few important calculations only. The values of rest are easily identifiable.

  • Cost of goods sold = materials + labor + heat, light & power
  • EBIT(i-t) = 70000*(1 – 0.40) = 70000*(0.60) = 42000
  • Total capital = common equity + long-term debt
  • Book value per share = Face value per share = (Common Equity / no. of shares)
  • Enterprise value = (Market Capitalisation + Debt) – Total Cash & Cash Equivalents
  • EBITDA = EBIT + Depreciation